# Detecting Out-of-distribution through the Lens of Neural Collapse

### v5

@@ Abstract
Efficient and versatile Out-of-Distribution (OOD) detection is essential for the safe deployment of AI yet remains challenging for existing algorithms. Inspired by Neural Collapse, we discover that features of in-distribution (ID) samples cluster closer to the weight vectors compared to features of OOD samples. In addition, we reveal that ID features tend to expand in space to structure a simplex Equiangular Tight Framework, which nicely explains the prevalent observation that ID features reside further from the origin than OOD features. Taking both insights from Neural Collapse into consideration, we propose to leverage feature proximity to weight vectors for OOD detection and further complement this perspective by using feature norms to filter OOD samples. Extensive experiments on off-the-shelf models demonstrate the efficiency and effectiveness of our method across diverse classification tasks and model architectures, enhancing the generalization capability of OOD detection.
@@ 1 Introduction
In this work, we focus on post-hoc methods, which address OOD detection independently of the training process. One line of prior work designs OOD scores over model output space (Djurisic et al., 2022; Hendrycks et al., 2019; Liang et al., 2018; Liu et al., 2020; Sun et al., 2021; Sun and Li, 2022) and another line of work focuses on the feature space, where OOD samples are observed to deviate from the clusters of ID samples (Lee et al., 2018; Mahalanobis, 2018; Sun et al., 2022; Tack et al., 2020). While existing research has made strides in OOD detection, they still face two major challenges: 1) maintaining detection effectiveness across different scenarios, and 2) ensuring computational efficiency for real-world deployment. For example, both output space and feature space methods suffer from performance discrepancy across different classification tasks, as shown in Table 1(a). Specifically, state-of-art algorithms on CIFAR-10 (Krizhevsky et al., 2009) OOD benchmarks perform suboptimally on ImageNet (Deng et al., 2009) OOD benchmarks, and vice versa. Such discrepancy is also observed across different architectures, as shown in Table 2. In addition, feature space methods, which rely on auxiliary models, raise efficiency concerns. For example, Lee et al. (2018) learns a Gaussian mixture model from training features and detects OOD based on Mahalanobis distance (Mahalanobis, 2018); Sun et al. (2022) records the training features and measures OOD-ness based on the k-th nearest neighbor distance to the training features. As shown in Liu and Qin (2023), such reliance on auxiliary models introduces additional cost, posing challenges for time-critical applications.
@@
Figure 1: Illustration of our framework inspired by Neural Collapse. Left: On the penultimate layer, ID samples cluster near the weight vectors (marked by stars) of their predicted classes while OOD samples reside separated, as shown by UMAP (McInnes et al., 2018). Middle: ID and OOD samples are separated by our proposed $\mathtt{pScore}$ (Equation 6), which measures feature proximity to weight vectors. Additionally, ID samples tend to reside further from the origin, illustrated with $\mathtt{L1}$ norms. Right: ID samples cluster near a simplex Equiangular Tight Framework, illustrated with black arrows denoting weight vectors. We detect OOD by thresholding on $\mathtt{pScore}$, selecting blue-shaded hypercones centered at weight vectors, with OOD samples outside these areas. Also, we filter OOD samples characterized by smaller feature norms. Left and Middle visualize a CIFAR-10 ResNet-18 classifier with OOD set SVHN. Right depicts our scheme on a three-class classifier with 2D penultimate space.
To this end, we aim to develop an efficient and versatile OOD detector. Specifically, we delve into the penultimate layer, i.e., the layer before the linear classification head, and revisit the observation that ID features tend to form clusters while OOD features reside apart. While this observation is well-established in prior literature (Lee et al., 2018; Sun et al., 2022; Tack et al., 2020), the underlying mechanism remains largely unexplained. To better understand and characterize the observation, we take insights from Neural Collapse (Papyan et al., 2020). Neural Collapse characterizes the interplay between the linear classification head and the penultimate layer features as training epochs go to infinity. Notably, the Neural Collapse phenomenons are observed to be prevalent across canonical classification tasks under diverse architectures (see Appendix E). One particular phenomenon that Neural Collapse reveals is that features of each class gradually converge towards a single point during training. This explains the prevalent clustering observed on off-the-shelf models as a precursor to the final convergence. Therefore, we leverage the landscape of Neural Collapse to understand:
@@
To address the question, we first demonstrate that as a deterministic effect of Neural Collapse, features of training samples will converge towards the weight vectors of the predicted class. Additionally, Neural Collapse reveals that training features also converge towards a simplex Equiangular Tight Framework (ETF) (Equation 1). The spatial structure of an ETF, illustrated in Figure 1 Right, corresponds to the maximum separation achievable in space by equiangular vectors, requiring the features to be sufficiently far from the origin.
Figure 2: Features of ID samples tend to cluster closer to the predicted class weight vectors, indicated by higher average cosine similarity (Equation 5) compared to OOD features. Evaluation across the training of a CIFAR-10 ResNet-18 classifier, with OOD set SVHN.
Such a landscape of Neural Collapse sheds light on the geometric structure of ID clusters. Specifically, for ID test samples, drawn from the same distribution as training samples, we anticipate a similar clustering behavior towards the weight vectors and towards an ETF. Conversely, OOD samples do not undergo the same training process, which enables the model to align features with weight vectors and to expand features to accommodate the spatial structure of ETF in Neural Collapse. Therefore, we do not expect the model to effectively align the weight vectors learned from ID features with unseen OOD features. Nor do we anticipate the model to posit OOD features far from the origin to structure an ETF. To validate our hypotheses, we trace a model’s training stages in Figure 2. We observe that ID samples consistently cluster closer to the weight vectors than OOD samples. This observation is reinforced in our UMAP (McInnes et al., 2018) visualization on an off-the-shelf CIFAR-10 classifier with ResNet-18 backbone in Figure 1 Left. Here, ID features cluster near predicted class weight vectors (marked by stars), while OOD features are distant. Combining our observation with the results in Zhu et al. (2021), which show that the weight vectors form an ETF in training, we conclude that ID features are driven to structure the ETF during training, whereas OOD features have no incentive to expand in space to form an ETF. Note that the lack of incentive for OOD features to expand in space explains the well-established observation (Huang et al., 2021; Park et al., 2023; Sun et al., 2022; Tack et al., 2020) that OOD features tend to reside closer to the origin.
@@
Based on our understanding, we design an efficient and versatile OOD detector. We first leverage feature proximity to the weight vectors to characterize ID clustering, bypassing auxiliary models and reducing the computational cost. Specifically, we define an angle-based proximity score as the norm of the projection of the weight vector of the predicted class onto the sample feature. As shown in Figure 1 Middle, our proximity score can effectively separate ID/OOD. A higher score indicates closer proximity and a lower chance of OOD-ness. Geometrically, thresholding on the score selects hyper-cones centered at the weight vector, as illustrated in Figure 1 Right. Notably, our proximity score effectively incorporates class-specific information and brings in performance benefits as well as efficiency gain. Complementing the proximity score’s contingency on ID clustering, we also consider feature distance to the origin. Specifically, ID features tend to reside further from the origin as they expand in space to form an ETF, whereas OOD features tend to reside near the origin, as illustrated by Figure 1 Right. Using the L1 norm as an example metric for distance to the origin, we observe that ID features can be separated from OOD features, as supported by Figure 1 Middle. Combining both aspects, we propose Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$).
@@
Notably, prior methods, e.g., KNN (Sun et al., 2022), focus on ID clustering but do not explicitly consider feature distance to the origin. Such approaches fall short in scenarios like ImageNet benchmarks but yield superior performance in CIFAR-10 benchmarks in Table 1(a). Conversely, methods such as Energy (Liu et al., 2020), Energy-based ASH (Djurisic et al., 2022), and, Energy-based Scale (Xu et al., 2023) inherently utilize feature distance to the origin by considering log-sum-exp of logits, yet largely overlook ID clustering. These approaches excel in some scenarios, e.g., ImageNet, but perform sub-optimally in others, e.g., CIFAR-10 (see Table 1(a)). Through the lens of Neural Collapse, we explain, connect, and complete prior methods under a holistic view, leading to improved efficiency and generalizability validated across diverse experiments.
@@
  • Understanding and Observation: By understanding ID clustering as a precursor to Neural Collapse, we novelly establish through experiments the significance of weight vectors in the clusters. We also explain from a spatial structure point of view the previous observation: ID features tend to reside further from the origin.
  • OOD Detector: We propose to leverage feature proximity to the weight vectors of predicted classes for OOD detection, which integrates class-specific information. Complementary to feature clustering, we also propose to detect OOD samples by thresholding the feature distance to the origin, which further enhances the method generalizability.
  • Experimental Analysis: We evaluate $\mathtt{NCI}$ across diverse classification tasks – CIFAR-10, CIFAR-100, ImageNet – and model architectures – ResNet, DenseNet (Huang et al., 2017), ViT (Dosovitskiy et al., 2020), Swin (Liu et al., 2022). Our $\mathtt{NCI}$ practically matches the latency of vanilla softmax-confidence detector while largely mitigating the performance discrepancy of existing methods across different classification tasks and model architectures.
2 Problem Setting
@@
In this work, we focus on detecting OOD samples from classes unseen during training, for which the classifiers cannot make meaningful predictions. The OOD detector $D:\mathcal{X}\rightarrow\{\text{ID},\text{OOD}\}$ is commonly constructed as: $D(\bm{x})=\begin{cases}\text{ID}&\text{if }s(\bm{x})\geq\tau\\ \text{OOD}&\text{if }s(\bm{x})<\tau\end{cases}$, where $s:\mathcal{X}\rightarrow\mathbb{R}$ is a score function of design and $\tau$ is the threshold. Considering the diverse application scenarios, an ideal OOD detector should be efficient and generalizable across classification tasks and model architectures. Thus, we propose a versatile and efficient OOD detector leveraging insights from Neural Collapse in this work.
@@ 3 OOD Detection through the Lens of Neural Collapse
In this section, we re-examine the observation in Lee et al. (2018) and Sun et al. (2022) that ID features tend to form clusters while OOD features deviate from the clusters. We understand the phenomenon as a precursor to the Neural Collapse (Papyan et al., 2020) convergence. And leveraging the landscape revealed in Neural Collapse, we examine:
Where do features of ID samples form clusters?
Through analytical and empirical study, we hypothesize and validate that (1) ID features tend to cluster closer to the weight vectors compared to OOD features; (2) ID clusters tend to reside further from the origin, as necessitated by their spatial structure. From our understanding, we develop a post-hoc OOD detector with enhanced efficiency and effectiveness.
3.1 Neural Collapse: Convergence of Training Features
Neural Collapse, first observed in Papyan et al. (2020), occurs on the penultimate layer across canonical classification settings during the Terminal Phase of Training (TPT) where training error vanishes and the training loss is trained towards zero. With $\bm{h}_{i,c}$ denoting the penultimate layer feature of the $i_{th}$ training sample with ground truth / predicted label $c$, Neural Collapse is framed in relation to
  • the feature global mean, $\bm{\mu}_{G}=\mathrm{Ave}_{i,c}\bm{h}_{i,c}$, where $\mathrm{Ave}$ is the average operation;
  • the feature class means, $\bm{\mu}_{c}=\mathrm{Ave}_{i}\bm{h}_{i,c},\ \forall c\in\mathcal{C}$;
  • the within-class covariance, $\bm{\Sigma}_{W}=\mathrm{Ave}_{i,c}(\bm{h}_{i,c}-\bm{\mu}_{c})(\bm{h}_{i,c}-\bm{\mu}_{c})^{T}$;
  • the between-class covariance, $\bm{\Sigma}_{B}=\mathrm{Ave}_{c}(\bm{\mu}_{c}-\bm{\mu}_{G})(\bm{\mu}_{c}-\bm{\mu}_{G})^{T}$;
  • the linear classification head, i.e. the last layer of the NN, $\arg\max_{c\in\mathcal{C}}\bm{w}_{c}^{T}\bm{h}+b_{c}$, where $\bm{w}_{c}$ and $b_{c}$ are parameters corresponding to class $c$.
Neural Collapse comprises four inter-related limiting behaviors:
@@ 3.1 Neural Collapse: Convergence of Training Features
We first remark on (NC2) that an ETF achieves the maximum separation possible for globally centered equiangular vectors (Papyan et al., 2020) and extends in space, as visualized in Figure 1 Right. Since training features converge towards an ETF, they need to have sufficient norms to accommodate the spatial arrangement.
We next build on (NC1) and (NC3) to demonstrate in the following that training features converge towards the weight vectors of the linear classification head, up to a scaling factor.
@@
$(\bm{h}_{i,c}-\bm{\mu}_{G})\rightarrow\lambda\bm{w}_{c}$ (2)
in the Terminal Phase of Training, where $\lambda=\frac{\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}}{\|\bm{w}_{c}\|_{2}}$.
@@
Considering that $(\bm{h}_{i,c}-\bm{\mu}_{c})(\bm{h}_{i,c}-\bm{\mu}_{c})^{T}$ is positive semi-definite for any $i$ and $c$. $\bm{\Sigma}_{W}\rightarrow\bm{0}$ thus implies $(\bm{h}_{i,c}-\bm{\mu}_{c})(\bm{h}_{i,c}-\bm{\mu}_{c})^{T}\rightarrow\bm{0}$ and $\bm{h}_{i,c}-\bm{\mu}_{c}\rightarrow\bm{0},\ \forall i,c$. With algebraic manipulations, we have
$\frac{\bm{h}_{i,c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}\rightarrow\bm{0},\ \forall i,c$ (3)
@@
$|\frac{\bm{h}_{i,c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{w}_{c}}{\|\bm{w}_{c}\|_{2}}|\leq|\frac{\bm{h}_{i,c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}|+|\frac{\bm{w}_{c}}{\|\bm{w}_{c}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu}_{G}}{\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}}|.$ (4)
@@
3.2 Geometric Structure of the Clusters of ID Features
While Neural Collapse reveals the within-class variability collapse (NC1), it also explains the clustering of features observed on general classifiers as a precursor to the collapse. We thus leverage the landscape of Neural Collapse revealed in Theorem 3.1 and (NC2) to examine the geometry of ID feature clusters. Since ID test samples are drawn from the same distribution as the training samples, we anticipate a similar pattern in their features. Specifically, we expect ID features to cluster towards the weight vectors of their predicted class during training. Additionally, we expect ID features to reside near a simplex Equiangular Tight Frame (ETF), thereby acquiring sufficient norm. Conversely, OOD samples are unseen during training and do not undergo the process of iterative adjustment, which drives the Neural Collapse phenomenon. Thus we expect the model to be less effective in aligning the OOD samples with weight vectors, placing OOD further from the weight vectors than ID features. Meanwhile, we do not expect the model to effectively align the OOD samples with an ETF.
In Figure 2, we validate our hypothesis across the training process of a CIFAR-10 classifier with ResNet-18 backbone. In Figure 2, we compute over the ID set (CIFAR-10) and OOD set (SVHN) the average cosine similarity between the centered feature $\bm{h}_{i}-\bm{\mu}_{G}$ and the weight vector $\bm{w}_{c}$ of the predicted class $c$, i.e.,
$Avg_{i}\ \ \frac{(\bm{h}_{i}-\bm{\mu}_{G})\cdot\bm{w}_{c}}{\|\bm{h}_{i}-\bm{\mu}_{G}\|_{2}\|\bm{w}_{c}\|_{2}}$ (5)
We observe that ID features have higher similarity scores and cluster closer to the weight vectors than OOD features. We further reinforce our observation in Figure 1 Left where we visualize ID features, OOD features, and weight vectors of a CIFAR-10 classifier with UMAP (McInnes et al., 2018). ID features are color-coded to align with the weight vectors (marked by stars) of their predicted classes, revealing a distinct clustering pattern near the weight vectors. Conversely, OOD features reside further away.
Additionally, we combine our observation with the results from Zhu et al. (2021), showing that the weight vectors form an ETF during training. Our observed proximity to the weight vectors thus also validates the clustering of ID features near an ETF and the divergence of OOD from this structure. The lack of structure and incentives to extend in space explains the relatively smaller norm of OOD features.

### v7

@@ Abstract
Out-of-Distribution (OOD) detection is critical for safe deployment; however, existing detectors often struggle to generalize across datasets of varying scales and model architectures, and some can incur high computational costs in real-world applications. Inspired by the phenomenon of Neural Collapse, we propose a versatile and efficient OOD detection method. Specifically, we re-characterize prior observations that in-distribution (ID) samples form clusters, demonstrating that, with appropriate centering, these clusters align closely with model weight vectors. Additionally, we reveal that ID features tend to expand into a simplex Equiangular Tight Frame, explaining the common observation that ID features are situated farther from the origin than OOD features. Incorporating both insights from Neural Collapse, our OOD detector leverages feature proximity to weight vectors and complements this approach by using feature norms to effectively filter out OOD samples. Extensive experiments on off-the-shelf models demonstrate the robustness of our OOD detector across diverse scenarios, mitigating generalization discrepancies and enhancing overall performance, with inference latency comparable to that of the basic softmax-confidence detector. Code is available at: https://github.com/litianliu/NCI-OOD.
@@ 1 Introduction
In this work, we focus on post-hoc methods, which address OOD detection independently of the training process. One line of prior work designs OOD scores over model output space [8, 13, 21, 24, 38, 37] and another line of work focuses on the feature space, where OOD samples are observed to deviate from clusters of ID samples [20, 27, 39, 40]. While existing research has made strides in OOD detection, they still face two major challenges: 1) maintaining detection effectiveness across different scenarios, and 2) ensuring computational efficiency for real-world deployment. For example, both output space and feature space methods suffer from performance discrepancy across different classification tasks, as shown in Table 1 (a). Specifically, strong algorithms on detecting OOD examples for CIFAR-10 [18] perform suboptimally in detecting OOD examples for ImageNet [5], and vice versa. No existing algorithm can simultaneously rank in the top three across two benchmarks, leading to sub-optimal average performance as shown in Table 1 (b). Furthermore, feature-space methods that rely on similarity measures to characterize ID clustering raise efficiency concerns. For example, [39] records training features and evaluates OOD-ness based on the k-th nearest neighbor distance to these features, while [23] demonstrates that dependency on such auxiliary models increases computational costs, posing challenges for time-sensitive applications. This highlights the need to move beyond feature-wise similarity and establish a more structured reference to characterize ID clustering efficiently.
Figure 1: Centered ID samples tend to cluster near the predicted class weight vectors, which are the last-layer weights of the corresponding class, as indicated by higher average cosine similarity (Equation 5) than OOD. This observation, inspired by the trend of Neural Collapse, emerges early in the training of this CIFAR-10 ResNet-18 classifier, with OOD set SVHN.
@@
Figure 2: Framework Illustration. Left: On the penultimate layer, the centered ID clusters reside near their predicted class weight vectors (marked by stars) while OOD samples reside separated, as shown by UMAP. Middle: ID and OOD samples are separated by $\mathtt{pScore}$ (Equation 6), which measures feature proximity to weight vectors. Also, ID samples tend to be further from the origin, illustrated with $\mathtt{L1}$ norms. Right: ID samples cluster near a simplex Equiangular Tight Framework, illustrated with black arrows denoting weight vectors. We detect OOD by thresholding on $\mathtt{pScore}$, selecting blue-shaded hypercones centered at weight vectors, with OOD samples outside these areas. We also filter OOD samples characterized by smaller feature norms. Left & Middle present a practical off-the-shelf CIFAR-10 ResNet-18 classifier with OOD set SVHN. Right depicts our scheme on a three-class classifier with 2D penultimate space.
To this end, we revisit the well-established observation that penultimate-layer ID features tend to form clusters, while OOD features reside apart [20, 39, 40], and take it a step further by asking:
@@
Leveraging Neural Collapse [32], we first examine the theoretical landscape with models trained beyond zero training error. We first show that, as a deterministic effect of Neural Collapse[32] penultimate-layer features of training samples converge toward the weight vectors of the predicted class—the last-layer weights of the corresponding class— after being centered by training feature mean. Additionally, Neural Collapse demonstrates that training features conform to a simplex Equiangular Tight Frame (ETF) (Equation 1), representing the maximum achievable separation between equiangular vectors. This spatial structure, illustrated in Figure 2 Right, requires features to be sufficiently distant from the origin.
Building on the insights from the theoretical convergence landscape, we propose two hypotheses about the geometric structure of ID clusters and OOD samples in practical models. Specifically, for ID test samples drawn from the same distribution as training samples, we expect a similar trend of clustering behavior toward the weight vectors and the ETF structure. In contrast, OOD samples, which are not part of the training process, lack the alignment with weight vectors and the spatial expansion necessary to form an ETF. Consequently, the model is unlikely to align unseen OOD features with weight vectors or place them far enough from the origin to exhibit ETF-like behavior. Although complete Neural Collapse requires strict conditions, such as prolonged training, its convergence trend and corresponding geometric patterns of ID features are consistently observed during training across diverse architectures and classification tasks (see Appendix 13 and [11]). Thus, our hypotheses regarding ID clusters and OOD samples are expected to hold without requiring complete Neural Collapse.
To validate the hypotheses, we trace a model’s training stages in Figure 1, where centered ID samples consistently cluster closer to the weight vectors than OOD samples. By epoch 50, such a pattern already enables effective OOD detection for SVHN samples  [^1] Our method (Section 3.3) achieves an AUROC of 94.44, outperforming the 91.27 softmax-confidence baseline (AUROC defined in Section 4).. This shows that the trend of Neural Collapse emerges early, validating the practical effectiveness of our hypothesis before full collapse in mature models. We reinforce this observation with a UMAP [28] visualization of an off-the-shelf CIFAR-10 classifier with a ResNet-18 backbone, shown in Figure 2 Left. ID features cluster near the predicted class weights (marked by stars), while OOD features remain distant. Additionally, as [53] shows that weight vectors form an ETF, our observations also support our second hypothesis: ID features are driven to structure the ETF, while OOD features lack the incentive to expand in space. This explains the well-established observation [40, 17, 39] that OOD features tend to reside closer to the origin, providing an alternative to the model confidence interpretation in [33].
@@
Based on our understanding, we design an efficient and versatile OOD detector. We first leverage feature proximity to the weight vectors to characterize ID clustering, effectively incorporating class-specific information and reducing the computational cost. Specifically, we define an angle-based proximity score as the norm of the projection of the weight vector of the predicted class onto the centered sample feature. As shown in Figure 2 Middle, our proximity score can effectively separate ID/OOD. A higher score indicates closer proximity and a lower chance of OOD-ness. Geometrically, thresholding on the score selects hyper-cones centered at the weight vector, as illustrated in Figure 2 Right. Complementing the proximity score’s contingency on ID clustering, we also consider feature distance to the origin. Specifically, ID features tend to reside further from the origin as they expand in space to form an ETF, whereas OOD features tend to reside near the origin, as illustrated by Figure 2 Right. Using the L1 norm as an example metric , we observe that ID features can be separated from OOD features, as supported by Figure 2 Middle. Combining both aspects, we propose Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$).
@@
Notably, prior methods, e.g., KNN [39], focus on ID clustering, without explicitly incorporating feature distance to the origin. Such approaches fall short in scenarios like ImageNet benchmarks but yield superior performance in CIFAR-10 benchmarks in Table 1(a). Conversely, methods such as Energy [24], Energy-based ASH [8], and, Energy-based Scale [47] inherently utilize feature distance to the origin by considering log-sum-exp of logits, yet largely overlook ID clustering. These approaches excel in scenarios like ImageNet, but perform sub-optimally in others, e.g., CIFAR-10. Through the lens of Neural Collapse, we explain, connect, and complete a wide range of prior methods under a holistic view, resulting in reduced latency and generalization discrepancies.
@@
  • Understanding and Observation: By analyzing ID clustering through the trend of Neural Collapse, we discover that penultimate-layer features of ID test examples cluster around the weight vectors of the predicted class after being centered by the training feature mean. Additionally, we explain that ID features tend to lie farther from the origin due to Neural Collapse, where they form a simplex Equiangular Tight Frame (ETF) structure. Notably, our findings hold without requiring full Neural Collapse convergence.
  • OOD Detector: We leverage feature proximity to the weight vectors of predicted classes for OOD detection, integrating class-specific information. Complementary to feature clustering, we propose to detect OOD samples by thresholding the feature distance to the origin.
  • Experimental Analysis: We evaluate $\mathtt{NCI}$ across diverse classification tasks (CIFAR-10, CIFAR-100, ImageNet) and model architectures (ResNet, DenseNet [15], ViT [9], Swin [26]). Rather than focusing on individual benchmarks, $\mathtt{NCI}$ reduces the generalization discrepancies and improves the overall effectiveness. We also show that incorporating weight vectors not only significantly improves upon prior clustering-based methods but also reduces latency to a minimal level, matching the basic softmax-confidence detector.
2 Problem Statement
@@
In this work, we focus on detecting OOD samples from classes unseen during training, for which the classifiers cannot make meaningful predictions. The OOD detector $D:\mathcal{X}\rightarrow\{\text{ID},\text{OOD}\}$ is commonly constructed as: $D(\bm{x})=\begin{cases}\text{ID}&\text{if }s(\bm{x})\geq\tau\\ \text{OOD}&\text{if }s(\bm{x})<\tau\end{cases}$, where $s:\mathcal{X}\rightarrow\mathbb{R}$ is a score function of design and $\tau$ is the threshold. Considering the diverse application scenarios, an ideal OOD detector should be efficient and generalizable. Thus, we leverage insights from Neural Collapse to achieve reduced computational costs and minimize generalization discrepancies.
@@ 3 OOD Detection through the Lens of Neural Collapse
In this section, we re-examine the observation in [20, 39] that ID features tend to form clusters while OOD features deviate from the clusters. We suggest that the clustering phenomenon can reflect the trend of the Neural Collapse [32] in practical models. From our understanding, we develop a post-hoc OOD detector with enhanced efficiency and effectiveness.
3.1 Neural Collapse: Convergence Landscape
Neural Collapse, first observed in [32], occurs on the penultimate layer across canonical classification settings. To formally introduce the concept, we use $\bm{h}_{c}^{i}$ to denote the penultimate layer feature of the $i_{th}$ training sample with ground truth / predicted label $c$, Neural Collapse is framed in relation to
  • feature global mean, $\bm{\mu}_{G}=\mathrm{Ave}_{i,c}\left(\bm{h}_{c}^{i}\right)$, where $\mathrm{Ave}\left(\cdot\right)$ is the average operation;
  • feature class means, $\bm{\mu}_{c}=\mathrm{Ave}_{i}\left(\bm{h}_{c}^{i}\right),\ \forall c\in\mathcal{C}$;
  • within-class covariance,
    $\bm{\Sigma}_{W}=\mathrm{Ave}_{i,c}\left((\bm{h}_{c}^{i}-\bm{\mu}_{c})(\bm{h}_{c}^{i}-\bm{\mu}_{c})^{T}\right);$
  • between-class covariance,
    $\bm{\Sigma}_{B}=\mathrm{Ave}_{c}\left((\bm{\mu}_{c}-\bm{\mu}_{G})(\bm{\mu}_{c}-\bm{\mu}_{G})^{T})\right);$
  • linear classification head, i.e. the last layer of the NN, $\arg\max_{c\in\mathcal{C}}\bm{w}_{c}^{T}\bm{h}+b_{c}$, where $\bm{w}_{c}$ and $b_{c}$ are parameters corresponding to class $c$.
Neural Collapse comprises four inter-related behaviors:
@@ 3.1 Neural Collapse: Convergence Landscape
We first remark on (NC2) that an ETF achieves the maximum separation possible for globally centered equiangular vectors [32] and extends in space, as visualized in Figure 2 Right. Since training features converge towards an ETF, they need to have sufficient norms to accommodate the spatial arrangement.
We next build on (NC1) and (NC3) to demonstrate that training features converge towards the weight vectors of the linear classification head, up to a scaling factor.
@@
$(\bm{h}_{c}^{i}-\bm{\mu}_{G})\rightarrow\lambda\bm{w}_{c},\vspace{-0.5mm}$ (2)
where $\lambda=\frac{\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}}{\|\bm{w}_{c}\|_{2}}$ in the Terminal Phase of Training.
@@
Considering that $(\bm{h}_{c}^{i}-\bm{\mu}_{c})(\bm{h}_{c}^{i}-\bm{\mu}_{c})^{T}$ is positive semi-definite for any $i$ and $c$. $\bm{\Sigma}_{W}\rightarrow\bm{0}$ thus implies $(\bm{h}_{c}^{i}-\bm{\mu}_{c})(\bm{h}_{c}^{i}-\bm{\mu}_{c})^{T}\rightarrow\bm{0}$ and $\bm{h}_{c}^{i}-\bm{\mu}_{c}\rightarrow\bm{0},\ \forall i,c$. With algebraic manipulations, we have
$\frac{\bm{h}_{c}^{i}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}\rightarrow\bm{0},\ \forall i,c$ (3)
@@
$|\frac{\bm{h}_{c}^{i}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{w}_{c}}{\|\bm{w}_{c}\|_{2}}|\leq$ (4)
$|\frac{\bm{h}_{c}^{i}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu_{G}}}{\|\bm{\mu}_{c}-\bm{\mu_{G}}\|_{2}}|+|\frac{\bm{w}_{c}}{\|\bm{w}_{c}\|_{2}}-\frac{\bm{\mu}_{c}-\bm{\mu}_{G}}{\|\bm{\mu}_{c}-\bm{\mu}_{G}\|_{2}}|.$
@@
3.2 Trend of Neural Collapse & Geometric Structure of the ID Clusters
While the complete collapse occurs during the Terminal Phase of Training (TPT) where training error vanishes and the training loss is trained towards zero, it is observed in [11] that the trend of Neural Collapse establishes in the early stages of training. We thus suggest that the clustering behavior of ID features observed in off-the-shelf models reflects a trend of Neural Collapse, corresponding to the within-class variability collapse (NC1). In light of this, we use the landscape of Neural Collapse revealed in Theorem 3.1 and (NC2) to examine the geometry of ID clusters.
Following our discussion in Section 1, we we hypothesize and validate with pre-trained models that (1) ID features tend to cluster closer to the weight vectors compared to OOD features; (2) ID clusters tend to reside further from the origin, as necessitated by their spatial structure. Particularly, Neural Collapse highlights the importance of centering for charaterizing ID clustering with weight vectors. We thus validate our hypothesis in a CIFAR-10 classifier with ResNet-18 backbone in Figure 1, computing over the ID set (CIFAR-10) and OOD set (SVHN) the average cosine similarity between the centered feature $\bm{h}^{i}-\bm{\mu}_{G}$ and the weight vector $\bm{w}_{c}$ of the predicted class $c$, i.e.,
$\mathrm{Ave}_{i}\left(\frac{(\bm{h}^{i}-\bm{\mu}_{G})\cdot\bm{w}_{c}}{\|\bm{h}^{i}-\bm{\mu}_{G}\|_{2}\|\bm{w}_{c}\|_{2}}\right)$ (5)